# Neurosymbolic Grounding for Compositional World Models

### v1

@@ Abstract
We introduce Cosmos, a framework for object-centric world modeling that is designed for compositional generalization (CG), i.e., high performance on unseen input scenes obtained through the composition of known visual “atoms." The central insight behind Cosmos is the use of a novel form of neurosymbolic grounding. Specifically, the framework introduces two new tools: (i) neurosymbolic scene encodings, which represent each entity in a scene using a real vector computed using a neural encoder, as well as a vector of composable symbols describing attributes of the entity, and (ii) a neurosymbolic attention mechanism that binds these entities to learned rules of interaction. Cosmos is end-to-end differentiable; also, unlike traditional neurosymbolic methods that require representations to be manually mapped to symbols, it computes an entity’s symbolic attributes using vision-language foundation models. Through an evaluation that considers two different forms of CG on an established blocks-pushing domain, we show that the framework establishes a new state-of-the-art for CG in world modeling.
@@ 1 Introduction
The discovery of world models — deep generative models that predict the outcome of an action in a scene made of interacting entities — is a central challenge in contemporary machine learning [12, 13]. As such models are naturally factorized by objects, methods for learning them [44, 10, 40, 21] commonly follow a modular, object-centric perspective. Given a scene represented as pixels, these methods first extract representations of the entities in the scene, then apply a transition network to model interactions between the entities. The entity extractor and the transition model form an end-to-end differentiable pipeline.
@@
Of particular interest in world modeling is the property of compositional generalization (CG), i.e., test-time generalization to scenes that are novel compositions of known visual “atoms". Recently, [44] gave a first approach to learning world models that compositionally generalize. Their method uses an action attention mechanism to bind actions to entities. The mechanism is designed to be equivariant to the replacement of objects in a scene by other objects, enabling CG.
@@
This paper continues the study of world models and CG. We note that CG is hard for purely neural methods, as they cannot easily learn encodings that can be decomposed into well-defined parts. Our approach, Cosmos, uses a novel form of neurosymbolic grounding to address this issue.
@@
The centerpiece idea in Cosmos is the notion of object-centric, neurosymbolic scene encodings. Like in prior modular approaches to world modeling, we extract a discrete set of entity representations from an input scene. However, each of these representations consistsd of: (i) a standard real vector representation constructed using a neural encoder, and (ii) a vector of symbolic attributes, capturing important properties — for example, shape, color, and orientation — of the entity.
@@
Like [10], we model transitions in the world as a collection of neural modules, each capturing a “rule” for pairwise interaction between entities. The mechanism for binding these rules to the entities is based on a novel neurosymbolic attention mechanism, in which the keys are symbolic, and the queries are neural. The symbolic keys are matched with the symbolic components of the entity encodings, enabling decisions such as “the $i$-th rule represents interactions between a black object and a circular object" (here, “black” and “circular” are symbolic attributes). The rules, the attention mechanism, and the entity extractor constitute an end-to-end differentiable pipeline.
@@
The CG abilities of this model stems from the compositional nature of the symbolic attributes. The symbols naturally capture “parts” of scenes. The neurosymbolic attention mechanism fires rules based on (soft, neurally represented) logical conditions over the symbols and can cover new compositions of the parts.
@@
A traditional issue with neurosymbolic methods for perception is that they need a human-provided mapping between perceptual inputs and their symbolic attributes [14, 32]. However, Cosmos automatically obtains this mapping from vision-language foundation models. Concretely, to compute the symbolic attributes of an object, we utilize Clip [27] to assign each object values from a set of known attributes.
@@
We compare Cosmos against a state-of-the-art baseline and an ablation on an established domain of moving 2D shapes [21]. Our evaluation considers two definitions of CG, illustrated in Figure 1. In one of these, we want the model to generalize to new scene compositions, i.e., input scenes comprising sets of entities that have been seen during training, but never together. The other definition, object composition, is new to this paper: here, we additionally need to accommodate shared dynamics between objects with shared attributes (e.g., the color red). Our results show that Cosmos outperforms the competing approaches at next-state prediction (Figure 1 visualizes a representative sample) and separability of the computed latent representations, as well as accuracy in a downstream planning task. Collectively, the results establish Cosmos as a state-of-the-art for CG world modeling.
@@
Figure 1: Overview of compositional world modeling. We depict examples from a 2D block pushing domain consisting of shapes that interact, where we can generate samples of different shapes and interactions. We aim to learn a model that generalizes to compositions not seen during training, such as scene composition (top) and object composition (bottom). Previous works [9] focus on scene composition, and struggle to generalize to harder compositional environments. Our approach Cosmos leverages object-centric, neurosymbolic scene encodings to compositionally generalize across settings containing different types of compositions.
@@
To summarize our contributions, we offer: (i) Cosmos, the first neurosymbolic, end-to-end differentiable approach — based on a combination of neurosymbolic scene encodings and neurosymbolic attention — to object-centric world modeling. (ii) a new way to use foundation models to automate symbol construction in neurosymbolic learning; (iii) an evaluation that shows Cosmos to produce significant empirical gains over the state-of-the-art for CG world modeling.
@@ 2 Problem Statement
We are interested in learning world models that compositionally generalize. World models arise naturally out of the formalism for Markov decision processes (MDPs). An MDP $\mathcal{M}$ is defined by a 5 tuple $(\mathcal{S},\mathcal{A},{T},{R},\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function ${T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\times\mathbb{R}_{\geq 0}$, immediate reward function ${R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ , and discount factor $\gamma\in[0,1]$. We make three additional constraints:
@@
Symbolic object relations: We assume that objects in the state space share attributes. An attribute is a set of unique symbols ${C_{p}}=\{{C_{p}^{1}},\dots C_{p}^{q}\}$. For instance, in the 2D block pushing domain (illustrated in Figure 1), each object has a “color” attribute that can take on values $C_{\texttt{color}}:=\{\texttt{red},\texttt{green},\dots\}$. An object can be composed of many such attributes that can be retrieved using an attribute projection function $\alpha_{\uparrow}:{S}_{i}\rightarrow\Lambda_{i}:=(\Lambda_{i}^{C_{1}};\dots;\Lambda_{i}^{C_{p}})$ where $(C_{1},\dots,C_{p})$ is a predefined, ordered list of attributes and $\Lambda_{i}^{C_{p}}$ selects the value in the $C_{p}$-th attribute that is most relevant to $S_{i}$. Note that $\alpha_{\uparrow}$ only depends on a single object and trivially generalizes to different compositions of objects [20].
Compositional Generalization: To formalize CG, we first formalize compositions. We assume that each state in our MDP can be decomposed using two sets of elements: compounds and atoms. Compounds are sets of elements that can be decomposed into smaller sets of elements, and atoms are sets of elements that cannot be decomposed further. For instance, in the block pushing domain (Figure 1), we can designate each unique object as an atom and designate the co-occurrence of a set of atoms in a scene as a compound. We use $\mathtt{A}$ to denote the elements designated to be the atoms and $\mathtt{C}$ to denote the compounds. The frequency distribution of the atoms is $\mathcal{F}_{\mathtt{A}}(\mathcal{D})$ and the frequency distribution of the compositions is $\mathcal{F}_{\mathtt{C}}(\mathcal{D})$. Given this, compositional generalization is expressed as a property of the train distribution $\mathcal{D}_{train}$ and of the test distribution $\mathcal{D}_{test}$ undergoing a distributional shift of the compounds, while the distribution of atoms remains the same.
Given the following assumptions, for any experience buffer $\mathcal{D}:=\{\{\langle I^{(t)},A^{(t)}\rangle\}_{t=1}^{T}\subseteq\mathcal{S}_{pixel}\times\mathcal{A}\}_{i=1}^{N}$, learning a world model boils down to learning the transition function $T$ for the MDP using the sequences of observations collected by $\mathcal{D}$. Specifically, for $I^{(i,t)},A^{(i,t)},I^{(i,t+1)}\in\mathcal{D}$, our objective function is
@@
Scene Composition: Figure 1 shows an instance of scene centric CG in the block pushing domain. Here, there exist $n=9$ unique objects, but only $k=3$ are allowed to co-occur in any given realized scene. Each unique object represents an atom, and the co-occurrence of a set of $k$ atoms in a scene represents the compound. Hence, there are a total $\binom{n}{k}$ possible compositions. The distribution of atoms in the train distribution and the test distribution does not change, while the distribution of compounds at train time and at test time are disjoint. So, $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{eval})$ is equivalent to $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{train})$ while $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})=\varnothing$.
Object Composition: Figure 1 shows an instance of object composition in the block pushing domain. Here, there are $n=9,k=3$ unique objects. A composition of objects occurs when two objects share attributes (for instance, here, the shared attributes are color and adjacency). Objects with shared attributes share dynamics. As a result, if one object experiences a force, others with the same attributes also undergo that force. Hence, assuming a single composition per scene, there are a total $\binom{n}{k}\binom{k}{2}$ possible compositions. As in the previous case, $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{eval})$ is equivalent to $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{train})$ while $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})=\varnothing$.
@@ 3 Method
Figure 2: Comparing world modeling frameworks between prior work [9] and Cosmos. Both modules start with entity extraction, to obtain neural object representations $\{S_{1},S_{2},S_{k}\}$ from the image (Section 3.1). While prior work uses this representation directly for the module selector, our work leverages a symbolic labeling module, which outputs a set of attributes $\Lambda$, to learn neurosymbolic representations (Section 3.2). We then perform action conditioning (Section 3.2.1) to keep track of corresponding actions, and update through a transition model (Section 3.3).
We approach neurosymbolic world modeling through object-centric neurosymbolic representations (Figure 2). Our method consists of: extracting object-centric encodings (Section 3.1), representing each object using neural and symbolic attributes (Section 3.2), and learning a transition model to update to next state (Section 3.3). Similar to previous works, we use slot based autoencoders to extract objects, but use symbolic grounding from foundation models alongside neural representations, which enables our model to achieve both scene and object compositionality. The full algorithm is presented in Algorithm 1, and the model is visualized in Figure 3.
Algorithm 1 Neurosymbolic world model for $k$-factorized state space. The input is an image $I$ of dimensions (C, H, W) = $(3\times 224\times 224)$ and a set of $k$-factorized actions of dimension $(k\times d_{\texttt{action}})$. The model is trained end-to-end except for the Slm modules, whose weights are held constant. There are three global variables: $T$, $\{C_{1},C_{2},C_{p}\}$ and $\{\vec{R}_{1},\vec{R}_{2},\vec{R}_{l}\}$. $T$ is the threshold of the number of repeated slot update steps, $\{C_{1},C_{2},C_{p}\}$ denotes text encodings for the closed vocabulary and $\{\vec{R}_{1},\vec{R}_{2},\vec{R}_{l}\}$ are learnable rule encodings.
1:function TransitionImage($I,[A_{1},A_{2},A_{k}]$)
2:   $\{S_{1},S_{2},S_{k}\}$leftarrow EntityExtractor($I$) $\triangleright$ $S_{i}$ dim: ($k,d_{\texttt{slot}}$)
@@
4:      $\{I_{1}^{\prime},I_{2}^{\prime},I_{k}^{\prime}\},\{M_{1}^{\prime},M_{2}^{\prime},M_{k}^{\prime}\}\leftarrow\textsc{SpatialDecoder}(\{S_{1},S_{2},S_{k}\})$triangleright $I_{i}$ dim: ($k$, C, H, W)
5:      $\{I_{1},I_{2},I_{k}\}\leftarrow\{M_{i}\cdot I_{i}^{\prime},\forall i\in[1,k]\}$
6:      $\{\Lambda_{1},\Lambda_{2},\Lambda_{k}\}\leftarrow\{\textsc{Slm}(I_{i},\{C_{1},C_{2},C_{p}\})|\forall i\in[1,k]\}$triangleright dim: ($k,p$)
7:      $\{\overline{\Lambda}_{1},\overline{\Lambda}_{2},\overline{\Lambda}_{k}\}\leftarrow\textsc{ActionAttn}(\{\Lambda_{1},\Lambda_{2},\Lambda_{k}\},[A_{1},A_{2},A_{k}])$
8:      ${p},{c},r\leftarrow\textsc{ModularRulenet}(\text{K=}\{\overline{\Lambda}_{1},\overline{\Lambda}_{2},\overline{\Lambda}_{k}\},\text{Q=}\{\vec{R}_{1},\vec{R}_{2},\vec{R}_{l}\})$
@@
10:   $\{I_{1}^{\prime},I_{2}^{\prime},I_{k}^{\prime}\},\{M_{1}^{\prime},M_{2}^{\prime},M_{k}^{\prime}\}\leftarrow\textsc{SpatialDecoder}(\{S_{1},S_{2},S_{k}\})$triangleright $I_{i}$ dim: ($k$, C, H, W)

### v2

@@ Abstract
We introduce Cosmos, a framework for object-centric world modeling that is designed for compositional generalization (CompGen), i.e., high performance on unseen input scenes obtained through the composition of known visual “atoms." The central insight behind Cosmos is the use of a novel form of neurosymbolic grounding. Specifically, the framework introduces two new tools: (i) neurosymbolic scene encodings, which represent each entity in a scene using a real vector computed using a neural encoder, as well as a vector of composable symbols describing attributes of the entity, and (ii) a neurosymbolic attention mechanism that binds these entities to learned rules of interaction. Cosmos is end-to-end differentiable; also, unlike traditional neurosymbolic methods that require representations to be manually mapped to symbols, it computes an entity’s symbolic attributes using vision-language foundation models. Through an evaluation that considers two different forms of CompGen on an established blocks-pushing domain, we show that the framework establishes a new state-of-the-art for CompGen in world modeling. Artifacts are available at https://trishullab.github.io/cosmos-web/.
@@ 1 Introduction
The discovery of world models — deep generative models that predict the outcome of an action in a scene made of interacting entities — is a central challenge in contemporary machine learning (Ha & Schmidhuber, 2018; Hafner et al., 2023). As such models are naturally factorized by objects, methods for learning them (Zhao et al., 2022; Goyal et al., 2021; Watters et al., 2019a; Kipf et al., 2019) commonly follow a modular, object-centric perspective. Given a scene represented as pixels, these methods first extract representations of the entities in the scene, then apply a transition network to model interactions between the entities. The entity extractor and the transition model form an end-to-end differentiable pipeline.
@@
Of particular interest in world modeling is the property of compositional generalization (CompGen), i.e., test-time generalization to scenes that are novel compositions of known visual “atoms". Recently, Zhao et al. (2022) gave a first approach to learning world models that compositionally generalize. Their method uses an action attention mechanism to bind actions to entities. The mechanism is equivariant to the replacement of objects in a scene by other objects, enabling CompGen.
@@
This paper continues the study of world models and compositional generalization. We note that such generalization is hard for purely neural methods, as they cannot easily learn encodings that can be decomposed into well-defined parts. Our approach, Cosmos, uses a novel form of neurosymbolic grounding to address this issue.
@@
The centerpiece idea in Cosmos is the notion of object-centric, neurosymbolic scene encodings. Like in prior modular approaches to world modeling, we extract a discrete set of entity representations from an input scene. However, each of these representations consists of: (i) a standard real vector representation constructed using a neural encoder, and (ii) a vector of symbolic attributes, capturing important properties — for example, shape, color, and orientation — of the entity.
@@
Like Goyal et al. (2021), we model transitions in the world as a collection of neural modules, each capturing a “rule” for pairwise interaction between entities. However, in contrast to prior work, we bind these rules to the entities using a novel neurosymbolic attention mechanism. In our version of such attention, the keys are symbolic and the queries are neural. The symbolic keys are matched with the symbolic components of the entity encodings, enabling decisions such as “the $i$-th rule represents interactions between a black object and a circular object" (here, “black” and “circular” are symbolic attributes). The rules, the attention mechanism, and the entity extractor constitute an end-to-end differentiable pipeline.
@@
The CompGen abilities of this model stems from the compositional nature of the symbolic attributes. The symbols naturally capture “parts” of scenes. The neurosymbolic attention mechanism fires rules based on (soft, neurally represented) logical conditions over the symbols and can cover new compositions of the parts.
@@
A traditional issue with neurosymbolic methods for perception is that they need a human-provided mapping between perceptual inputs and their symbolic attributes (Harnad, 1990; Tang & Ellis, 2023). However, Cosmos automatically obtains this mapping from vision-language foundation models. Concretely, to compute the symbolic attributes of an object, we utilize Clip (Radford et al., 2021) to assign each object values from a set of known attributes.
@@
We compare Cosmos against a state-of-the-art baseline and an ablation on an established domain of moving 2D shapes (Kipf et al., 2019). Our evaluation considers two definitions of CompGen, illustrated in Figure 1. In one of these, we want the model to generalize to new entity compositions, i.e., input scenes comprising sets of entities that have been seen during training, but never together. The other definition, relational composition, is new to this paper: here, we additionally need to accommodate shared dynamics between objects with shared attributes (e.g., the color red). Our results show that Cosmos outperforms the competing approaches at next-state prediction (Figure 1 visualizes a representative sample) and separability of the computed latent representations, as well as accuracy in a downstream planning task. Collectively, the results establish Cosmos as a state-of-the-art for CompGen world modeling.
@@
Figure 1: Overview of compositional world modeling. We depict examples from a 2D block pushing domain consisting of shapes that interact, where we can generate samples of different shapes and interactions. We aim to learn a model that generalizes to compositions not seen during training, such as entity composition (top) and relational composition (bottom). Previous works (Goyal et al., 2021) focus on entity composition, and struggle to generalize to harder compositional environments. Our approach Cosmos leverages object-centric, neurosymbolic scene encodings to compositionally generalize across settings containing different types of compositions.
To summarize our contributions, we offer:
@@
  • Cosmos, the first differentiable neurosymbolic approach — based on a combination of neurosymbolic scene encodings and neurosymbolic attention — to object-centric world modeling.
  • a new way to use foundation models to automate symbol construction in neurosymbolic learning;
  • an evaluation that shows Cosmos to produce significant empirical gains over the state-of-the-art for compositionally generalizable world modeling.
@@ 2 Problem Statement
We are interested in learning world models that compositionally generalize. World models arise naturally out of the formalism for Markov decision processes (MDPs). An MDP $\mathcal{M}$ is a tuple $(\mathcal{S},\mathcal{A},{T},{R},\gamma)$ with states $\mathcal{S}$, actions $\mathcal{A}$, transition function ${T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\times\mathbb{R}_{\geq 0}$, reward function ${R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ , and discount factor $\gamma\in[0,1]$. We make three additional constraints:
@@
Symbolic object relations: We assume that objects in the state space share attributes. An attribute is a set of unique symbols ${C_{p}}=\{{C_{p}^{1}},\dots C_{p}^{q}\}$. For instance, in the 2D block pushing domain (illustrated in Figure 1), each object has a “color” attribute that can take on values $C_{\texttt{color}}:=\{\texttt{red},\texttt{green},\dots\}$. An object can be composed of many such attributes that can be retrieved using an attribute projection function $\alpha_{\uparrow}:{S}_{i}\rightarrow\Lambda_{i}:=(\Lambda_{i}^{C_{1}};\dots;\Lambda_{i}^{C_{p}})$ where $(C_{1},\dots,C_{p})$ is a predefined, ordered list of attributes and $\Lambda_{i}^{C_{p}}$ selects the value in the $C_{p}$-th attribute that is most relevant to $S_{i}$. Note that $\alpha_{\uparrow}$ only depends on a single object and trivially generalizes to different compositions of objects (Keysers et al., 2020).
[…]
Given the following assumptions, for any experience buffer $\mathcal{D}:=\{\{\langle I^{(t)},A^{(t)}\rangle\}_{t=1}^{T}\subseteq\mathcal{S}_{pixel}\times\mathcal{A}\}_{i=1}^{N}$, learning a world model boils down to learning the transition function $T$ for the MDP using the sequences of observations collected by $\mathcal{D}$. Specifically, for $S^{(i,t)}_{1\dots k}\in I^{(i,t)},A^{(i,t)},I^{(i,t+1)}\in\mathcal{D}$, our objective function is
@@
Entity Composition: Figure 1 shows an instance of entity centric CompGen in the block pushing domain. Here, there exist $n=9$ unique objects, but only $k=3$ are allowed to co-occur in any given realized scene. Each unique object represents an atom, and the co-occurrence of a set of $k$ atoms in a scene represents the compound. Hence, there are a total $\binom{n}{k}$ possible compositions. The distribution of atoms in the train distribution and the test distribution does not change, while the distribution of compounds at train time and at test time are disjoint. So, $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{eval})$ is equivalent to $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{train})$ while $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{train})=\varnothing$.
Relational Composition: Figure 1 shows an instance of relational composition in the block pushing domain. Here, there are $n=9,k=3$ unique objects. A composition of objects occurs when two objects share attributes (for instance, here, the shared attributes are color and adjacency). Objects with shared attributes share dynamics. As a result, if one object experiences a force, others with the same attributes also undergo that force. Hence, assuming a single composition per scene, there are a total $\binom{n}{k}\binom{k}{2}$ possible compositions. As in the previous case, $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{eval})$ is equivalent to $\mathcal{F}_{\mathtt{A}}(\mathcal{D}_{train})$ while $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{train})=\varnothing$.
@@ 3 Method
Figure 2: Comparing world modeling frameworks between prior work (Goyal et al., 2021) and Cosmos. Both modules start with entity extraction, to obtain neural object representations $\{S_{1},\dots S_{k}\}$ from the image (Section 3.1). While prior work uses this representation directly for the module selector, our work leverages a symbolic labeling module, which outputs a set of attributes $\Lambda$, to learn neurosymbolic representations (Section 3.2). We then perform action conditioning (Section 3.2) to keep track of corresponding actions, and update through a transition model (Section 3.3).
We approach neurosymbolic world modeling through object-centric neurosymbolic representations (Figure 2). Our method consists of: extracting object-centric encodings (Section 3.1), representing each object using neural and symbolic attributes (Section 3.2), and learning a transition model to update to next state (Section 3.3). Similar to previous works, we use slot based autoencoders to extract objects, but use symbolic grounding from foundation models alongside neural representations, which enables our model to achieve both entity and relational compositionality. The full algorithm is presented in Algorithm 1, and the model is visualized in Figure 3. We will make use of the example in Figure 3 to motivate our algorithm.
Algorithm 1 Neurosymbolic world model for $k$-factorized state space. The input is an image $I$ of dimensions (C, H, W) = $(3\times 224\times 224)$ and a set of $k$-factorized actions of dimension $(k\times d_{\texttt{action}})$. The model is trained end-to-end except for the Slm modules, whose weights are held constant. There are three global variables: $T$, $\{C_{1},\dots C_{p}\}$ and $\{\vec{R}_{1},\dots\vec{R}_{l}\}$. $T$ is the threshold of the number of repeated slot update steps, $\{C_{1},\dots C_{p}\}$ denotes text encodings for the closed vocabulary and $\{\vec{R}_{1},\dots\vec{R}_{l}\}$ are learnable rule encodings.
1:function TransitionImage($I,[A_{1},\dots\dots A_{k}]$)
2:   $\{S_{1},\dots S_{k}\}$leftarrow EntityExtractor($I$) $\triangleright$ $S_{i}$ dim: ($k,d_{\texttt{slot}}$)
@@
4:      $\{I_{1}^{\prime},\dots I_{k}^{\prime}\},\{M_{1}^{\prime},\dots M_{k}^{\prime}\}\leftarrow\textsc{SpatialDecoder}(\{S_{1},\dots S_{k}\})$triangleright $I_{i}^{\prime}$ dim: (C, H, W)
5:      $\{I_{1},\dots I_{k}\}\leftarrow\{M_{i}^{\prime}\cdot I_{i}^{\prime},\forall i\in[1,k]\}$
6:      $\{\Lambda_{1},\dots\Lambda_{k}\}\leftarrow\{\textsc{Slm}(I_{i},\{C_{1},\dots C_{p}\})|\forall i\in[1,k]\}$triangleright dim: ($k,p$)
7:      $\{\overline{\Lambda}_{1},\dots\overline{\Lambda}_{k}\}\leftarrow\textsc{ActionAttn}(\{\Lambda_{1},\dots\Lambda_{k}\},[A_{1},\dots A_{k}])$
8:      ${p},{c},r\leftarrow\textsc{ModularRulenet}(\text{K=}\{\overline{\Lambda}_{1},\dots\overline{\Lambda}_{k}\},\text{Q=}\{\vec{R}_{1},\dots\vec{R}_{l}\})$
@@
10:   $\{I_{1}^{\prime},\dots I_{k}^{\prime}\},\{M_{1}^{\prime},\dots M_{k}^{\prime}\}\leftarrow\textsc{SpatialDecoder}(\{S_{1},\dots S_{k}\})$triangleright $I_{i}^{\prime}$ dim: (C, H, W)